Please write clear

Solve the following problem using Karush-Kuhn-Tucker necessary conditions:

Maximize f(X) = 8x1 + 4x2 + x1x2 - x12 - x22 subject to:

g1(X): 2x1 + 3x2 ≤ 24, g2(X): -5x1 + 12x2 ≤ 24, g3(X): x2 ≤ 5.
Answer #1

Solh Map flor) = 8 ubjed to 804 + Alla +242-re, ² - ² 372 224 -) g, a) 224 t 54, 12.12s24 Izm <5: 9₃ lm) no, n2zo. The LagranCase let 2, 20.= d = ² from N2-224 +8 20 = -224 + s = -8 2.3Um2 = 53 no 202 +4=0 24532_4= -372 = -16 2G 3 32.12 3 pia 20 froCase 4, 20, dato 20 froma o-me- eurot lus- -O frono 8 + + an, -3d2=0 324 Un2-829 -12, A *14, 222-1222 = 0 = +222 +2 -1 dae